# LARA: Linguistic-Adaptive Retrieval-Augmented LLMs for Multi-Turn Intent Classification

### v1

@@ 5 Experiments
5.1 Dataset
The datasets used in this work consist of user queries in the local languages of eight markets: Brazil, Indonesia, Malaysia, Philippines, Singapore, Thailand, Taiwan, and Vietnam. The queries are related to the E-commerce domain.
Market Lang. Intents Train(Single-Turn) Test(Multi-Turn)
@@ 5.1 Dataset
Table 1: The major languages, number of intents, and the number of samples in each market.
Table 1 shows the number of samples we have in each dataset. All the data are collected through the manual annotation by local CS teams of each market. We have the single-turn training data available in abundance over the course of business operation after years. These single-turn samples will serve as the demonstration pool for in-context learning. To evaluate the effectiveness of our methods, we also have the CS teams to manually annotate some real multi-turn online sessions to serve as the test set. Each session queries $\mathcal{Q}$ will only have the last query $q_{n}$ labelled.
5.2 Metrics
We evaluate the accuracy of the methods based only on the label of the last query $q_{n}$ in each conversation session $\mathcal{Q}$. Other metrics which consider class imbalance are not used as the sampled sessions are expected to reflect the online traffic of each intent, thus better simulates the true online performance.
5.3 Baselines
To the best of our knowledge, there is no any existing work that directly addresses the challenge of multi-turn intent recognition with a large number of classes. This is a challenging task due to the lack of labelled data.
In our work, we present two intuitive and realistic approaches as baselines:
5.3.1 Naive concatenation
All queries in a single session $\mathcal{Q}$ are concatenated using the $\circ$ operation mentioned above, and the concatenation result is fed into the single-turn model $\mathcal{M}_{c}$ for inference.
5.3.2 Selective concatenation
In this approach, only one query from $\mathcal{C}_{q}$ is selected to be concatenated with $q_{n}$. The intuition is that not all history queries are helpful in understanding the last query, and the excessive use of them might introduce unwanted noise. A concatenation decision model is trained to select the most suitable history query. Depending on the model confidence, there might be cases where no expansion is needed at all.
5.4 Implementation Details
The traditional single-turn model, the retriever, and the concatenation decision model used are using backbone initialized with $\Phi_{XLMR}$, a multi-lingual domain specific XLM-RoBERTa-base model continued to be pre-trained with contrastive learning. We use AdamW to finetune the backbone and all other modules with a learning rate of 5e-6 and 1e-3, respectively. In LARA, the LLM used is vicuna-13b-v1.5 on Hugging Face with 13B parameters. All test are run on a single Nvidia V100 GPU card with a 32GB of GPU memory. The number of demonstrations $K$ retrieved for each intent is set at 10 in this experiment. Due to GPU memory constraint, the total number of tokens the in-context learning demonstrations can make up to are limited to 2300 tokens. If exceeded, the number of demonstrations in each candidate intent are pruned equally starting with the ones with the lowest cosine similarity scores to $q_{all}$. During inference time, if the generated intent doesn’t match any of the provided options, the intent of $\mathcal{M}_{c}$ on $q_{n}$ will be considered as the final result.
6 Results and Discussions
Table 2 compares the performance of baselines and LARA. On average, LARA achieves better results than the baselines in any of the prompt variants used. Naive concatenation is not always more effective than Selective concatenation, showing that naively including all history queries will introduce noises which in turn jeopardizes the performance. However, pseudo-labelling the dataset used to train the concatenation decision model will need to be carefully carried out, and despite the extra steps, it will not necessarily be more effective than the naive method. LARA, on the other hand, can achieve good results on most dataset without any complex pseudo-labelling process. This also highlights the linguistic-adaptivity of the method on broad languages. The only market that it doesn’t outperform the baselines is ID, which most probably can be attributed to the language ability of open-sourced LLMs in handling the local slang and abbreviations in casual conversation. After all, the backbone model used in baselines are pre-trained directly on the in-domain chat log data, while the LLM models are used out-of-the-box.
Model Prompt BR ID MY PH SG TH TW VN avg
Replacing the label names with non-related symbols in $\mathcal{P}_{symbolic}$ significantly hurts the performance of in-context learning. On the other hand, minimal changes to label names in $\mathcal{P}_{prepend}$ does not heavily impact the performance. In turn, the inference time is improved by 77%, from 0.75it/s to 1.32it/s on a single V100 card using Hugging Face python library. Interestingly, the model also stopped generating labels which cannot be matched with the options provided in demonstrations, while previously the rate is on average 1.6% using $\mathcal{P}$. Finally, we also tried a new prompt $\mathcal{P}_{formatted}$ based on $\mathcal{P}_{prepend}$. Only a very slight change to the context format is done, but it can outperform the other prompt variants in all dataset, suggesting that giving $\mathcal{C}_{\mathcal{Q}}$ a closer format to $\mathcal{E}$ and the targeted $q_{n}$ will be more beneficial in the context utilization. Besides, this also hints that the prompt could also be worked on more in the future as it is not extensively tuned in this work.
7 Ablation Studies
To validate our motivation and model design, we ablate our model components. The comparison is made on the original $\mathcal{P}$ prompt variant.
Figure 3: Ablation on different components of LARA. The last row shows the standard deviations of performance ‘w/o Retrieval’ over 10 runs. The size of the demonstration pool and the number of intents in each dataset are also included here for the ease of reference.
7.1 Traditional Single-turn Model
We tried not to use $\mathcal{M}_{c}$ to narrow down the intent candidates before retrieving the demonstrations. All demonstrations are directly retrieved based on their cosine similarity to $q_{all}$. As the original number of retrieved demonstrations for each $\mathcal{Q}$ is dynamic according to number of queries $n$, to ensure fairness, the number of retrieved demonstrations in this variant will also be $K\times n$. From Fig LABEL:fig:enter-label, we see that the quality of in-context learning is adversely impacted when the number of intents included in demonstrations is too high.
7.2 Retrieval Component
Demonstration selection could also have significant impact on the performance. Thus, we also tried to remove the retrieval component, and randomly sample the demonstrations for each intent. The results are reported with 10 runs on the random sampling. Based on Table LABEL:fig:enter-label, the overall performance will worse if there is no retrieval component, specifically when there are big demonstration pool or high number of intents. Thus, if there is no one good strategy in pruning the demonstration pool, the easier way is to retrieval demonstrations based on their similarity to the task input.
8 Conclusion
This paper introduced LARA, a framework that leverages Linguistic-Adaptive Retrieval-Augmentation to address multi-turn intent classification challenges through zero-shot settings across multiple languages. Unlike other supervised Fine-Tuning (SFT) models, which require a multi-turn dialogue set that is hard to collect. Our method only requires a single-turn training set to train a conventional XLM model. It then combines it with an innovative in-context retrieval augmentation for multi-turn intent classification. LARA demonstrated a notable improvement in accuracy and efficiency, marking a significant advancement in the field of conversational AI.

### v2

@@ 5 Experiments
5.1 Experimental Setup
Dataset. The dataset is obtained from the conversation history of a large e-commerce company. It consists of user queries in the local languages of eight markets: Brazil, Indonesia, Malaysia, Philippines, Singapore, Thailand, Taiwan, and Vietnam. All the labelled data are collected through manual annotation by local customer service teams of each market. The samples with consistency labels from 3 taggers are selected to ensure the annotation quality. More details are in B
Metrics. We evaluate the accuracy of the methods based only on the label of the last query $q_{n}$ in each conversation session $\mathcal{Q}$. Other metrics which consider class imbalance are not used as the sampled sessions are expected to reflect the online traffic of each intent, thus more accurately simulating true online performance.
5.2 Baselines
Current multi-turn models are trained with multi-turn datasets, but our methods did not require such data. For a fair comparison, we adopt a state-of-the-art single-turn model mentioned in 4.1 with two types of concatenation approaches as baselines:
5.2.1 Single-turn approach
Inference on only the last utterance of a session using the single-turn model $\mathcal{M}_{c}$, all previous contexts are ignored.
5.2.2 Naive concatenation
All queries in a single session $\mathcal{Q}$ are concatenated using comma, and the concatenation result is fed into the single-turn model $\mathcal{M}_{c}$, which is enhanced by HTC approach mentioned above, for inference. HTC methods which optimally utilise the overall label hierarchy information often outperform the methods which simply disregard the structureRojas et al. (2020).
5.2.3 Selective concatenation
In this approach, only one query from $\mathcal{C}_{q}$ is selected to be concatenated with $q_{n}$. The intuition is that not all history queries are helpful in understanding the last query, and the excessive use of them might introduce unwanted noise. A concatenation decision model is trained to select the most appropriate historical query. Depending on the model confidence, there might be cases where no expansion is needed at all. The concatenation result is then also fed into the single-turn model $\mathcal{M}_{c}$ for inference.
6 Results and Discussions
Table 2 compares the performance of baselines and LARA on our multi-turn dataset. The single-turn approach has the worst performance due to the lack of context from history queries. The single-turn model with Naive concatenation is lower than Selective concatenation by 0.89% on average, showing that naively including all history queries will introduce noises, which in turn jeopardizes the performance. However, pseudo-labelling the dataset used to train the concatenation decision model will need to be carefully carried out, and despite the extra steps, it will not necessarily be more effective than the naive method.
LARA, on the other hand, with prompt $\mathcal{P}_{formatted}$, achieves the best results on most markets without any multi-turn training data. On average, it improved accuracy by 3.67% compared with Selective concatenation. Especially on the non-English markets, it also improved by 2.96%, 4.18% and 4.00% on TH, VN and BR separately. This highlights the linguistic-adaptivity of the method on broad languages. The only market that does not outperform the baselines is ID, which most probably can be attributed to the language ability of open-sourced LLMs in handling the local slang and abbreviations in casual conversation. After all, the backbone model used in baselines is pre-trained directly on the in-domain chat log data, while the LLM models are used out of the box.
Model Prompt BR ID MY PH SG TH TW VN avg
Single-turn Approach - 30.98% 52.14% 56.81% 40.21% 51.13% 52.99% 58.07% 65.90% 53.76%
Replacing the label names with non-related symbols in $\mathcal{P}_{symbolic}$ significantly hurts the performance of in-context learning. On the other hand, minimal changes to label names in $\mathcal{P}_{prepend}$ does not heavily impact the performance. In turn, the inference time is improved by 77%, from 0.75it/s to 1.32it/s on a single V100 card using Hugging Face python library. Interestingly, the model also stopped generating labels which cannot be matched with the options provided in demonstrations, while previously the rate is on average 1.6% using $\mathcal{P}$. Finally, we also tried a new prompt $\mathcal{P}_{formatted}$ based on $\mathcal{P}_{prepend}$. Only a very slight change to the context format is done, but it can outperform the other prompt variants in all datasets, suggesting that giving $\mathcal{C}_{\mathcal{Q}}$ a closer format to $\mathcal{E}$ and the targeted $q_{n}$ will be more beneficial in the context utilization. Besides, this also hints that the prompt could also be worked on more in the future as it is not extensively tuned in this work.
Encoder BR ID MY PH SG TH TW VN avg
Our own encoder 55.65% 62.88% 64.71% 55.03% 65.40% 61.95% 66.86% 78.10% 64.64%
all-MiniLM-L12-v2 54.57% 61.05% 64.36% 53.97% 65.13% 59.56% 68.27% 75.81% 63.63%
all-mpnet-base-v2 55.91% 60.79% 63.30% 54.50% 64.31% 58.57% 67.42% 76.95% 63.24%
paraphrase-multilingual
-MiniLM-L12-v2 54.03% 62.01% 64.42% 57.14% 64.31% 62.35% 69.12% 76.95% 64.25%
paraphrase-multilingual
-mpnet-base-v2 55.38% 61.57% 65.42% 55.56% 64.99% 59.96% 68.56% 76.95% 64.29%
Table 3: Performance of LARA using different encoders for in-context demonstration mining. The best performance for each dataset is in bold, while the second best is underlined.
7 Ablation Studies
To validate our motivation and model design, we ablate single-turn model $\mathcal{M}_{c}$ in candidate intent selection and retrievers in ICL retrieval. The comparison is made on the original $\mathcal{P}$ prompt variant.
7.1 The necessity of model $\mathcal{M}_{c}$
$\mathcal{M}_{c}$ is used to recognise the intent candidates before ICL retrieval. If so, all demonstrations are directly retrieved based on their cosine similarity to $q_{all}$ and the quality of in-context learning is adversely impacted. The accuracy on all markets dramatically dropped except PH, which only dropped by 0.53% due to the least number of intents. The average score across all markets dropped from 64.18% to 53.99%. For instance, "refund timeline" and "refund timeline for cancelled order" could be confusing to retrieval-based models, while classification models trained on each market dataset can discern them better.
Figure 3: Ablation on different components of LARA.
7.2 The role of retrievers in ICL retrieval
Demonstration selection via retrievers may have an impact on the performance. Thus, we remove all retrievers and randomly sample the demonstrations for each intent. The results are reported with 10 runs on the random sampling. As shown in Figure 3, the overall performance decreased by 0.71% without retrievers, highlighting the importance of selecting demonstrations which are more similar to test queries. ID and VN, with demonstration pool two to three times bigger than others (refer to Appendix B for pool size), are affected the most because the chance of selecting samples which are not so similar is higher.
7.3 Quality of Embedding Model
We also experimented with open-source, multilingual semantic similarity models provided by SentenceTransformers Reimers and Gurevych (2019) for in-context demonstration mining. These models are easily accessible to the public and cover a wide range of languages. Due to the training method, the similarity metric used is still cosine similarity.
The multilingual models selected in this experiment have lower or roughly the same number of parameters as our own XLM-RoBERTa-base model used in the paper. They are
  • Models with all-rounded abilities which have the best average performance reported by the authors: all-MiniLM-L12-v2 and all-mpnet-base-v2
  • Models for paraphrase mining as it is similar to our task of comparing multi-turn utterances to single-turn utterances: paraphrase-multilingual-MiniLM-L12-v2 and paraphrase-multilingual-mpnet-base-v2
All comparisons are done using the same prompt, $\mathcal{P}_{formatted}$. From the table 3, the best open-source model of selections (from row 2 - 5) is overall only 0.35% behind our own model, which means that the quality of the embedding model does not matter much and can easily be replaced by open-source models. That said, we would recommend paraphrase mining models over the general purpose ones as they are more suited for the scenario. Furthermore, if resource is a concern, smaller models like MiniLM-L12 can be selected with 3x speed up compared to mpnet-base, all while maintaining the same overall quality.
8 Limitation
LARA does not currently address the detection of out-of-domain utterances, a critical aspect for online dialogue systems. Future research is necessary to explore methods for incorporating this capability and to assess their feasibility. Furthermore, the resilience of the method to user intent shifts has not been examined. Additionally, the multi-component architecture, which integrates text classification, retrieval, and ICL, adds to the implementation complexity. In Appendix D, we propose a lightweight solution that is more suitable for applications with limited deployment resources.
9 Conclusion
This paper introduced LARA, a framework that leverages Linguistic-Adaptive Retrieval-Augmentation to address multi-turn intent classification challenges through zero-shot settings across multiple languages. Unlike other supervised Fine-Tuning (SFT) models, which require a hard-to-collect multi-turn dialogue set, our method requires only a single-turn training set to train a conventional model, combining it with an innovative in-context retrieval augmentation for multi-turn intent classification. LARA substantially improved user satisfaction by 1% over multi-turn sessions, reduced the transfer-to-agent rate by 0.5% and saved the cost of hundreds of agents, which is a key business metric in our industrial application.
Acknowledgments. This research is supported in part by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Award No. MOE-T2EP20123-0015). Any opinions, findings and conclusions, or recommendations expressed in this material are those of the authors and do not reflect the views of the Ministry of Education, Singapore.
Appendix A The details of single-turn model ($\mathcal{M}_{c}$)
A text classification model is trained on the annotated single-turn dataset $\mathcal{D}$. The model is an ensemble of a simple label-attention model and a state-of-the-art global HTC approach. The label-attention model exploits local information per layer of the taxonomy by having separate classifier head for each intent taxonomy layer, whereas the global approach addresses the task with a single model for all the classes and levels. In our implementation, both approaches are trained as a single network and back propagation is performed on the ensembled output.
The approaches share the same encoder and sentence representation. Given a query $q$, we adopt the [CLS] token embedding from XLM-RoBERTa-base model with weight $\Phi_{XLMR}$ as the text representation $H$. Formally,
$H=\Phi_{XLMR}(q)^{[CLS]}\in\mathbb{R}^{d}$
where $d$ is the hidden dimension. $\Phi_{XLMR}$ had been further pretrained with our in-domain corpus to give meaningful representation to [CLS] token.
In the label-attention model, we have one classifier head for each intent layer. Each of the classifier heads has one hidden linear layer to obtain the layer intermediate output $L_{l}$, which encodes the layer information. This layer information will be utilised in the input of of the next layer classifier head.
$L_{l}=\begin{cases}HW^{1}_{l}+b^{1}_{l},&\text{if }l=1,\\ (H\oplus L_{l-1})W^{1}_{l}+b^{1}_{l},&\text{if }l>1,\end{cases}$
where $W^{1}_{l}\in\mathbb{R}^{d\times d}\text{ for }l=1\text{ and }W^{1}_{l}\in\mathbb{R}^{2d\times d}\text{ for }l>1$. $b^{1}_{l}\in\mathbb{R}^{d}$, $l$ is the layer number, $\oplus$ denotes tensor concatenation. Finally, we obtain the local logits $H_{local}^{l}$ for each layer classes by using another linear layer
$H_{local}^{l}=L_{l}\cdot W^{2}_{l}+b^{2}_{l},W^{2}_{l}\in\mathbb{R}^{d\times|\mathcal{I}_{l}|},b^{2}_{l}\in\mathbb{R}^{|\mathcal{I}_{l}|}$
where $|\mathcal{I}_{l}|$ is the number of classes in the layer.
However, the label-attention model used is not aware of the overall hierarchical structure. Therefore, we ensemble it with another method. We refer to HiTIN Zhu et al. (2023) for the implementation of state-of-the-art HTC global approach. In this method, a tree network is constructed based on the simplified original taxonomy structure, and the messages are propagated bottom-up in an isomorphism manner, which complements the label-attention model used. The embedding for leaf nodes are obtained by broadcasting the text representation $H$. After the tree isomorphism network propagation, all embedding from all layers are aggregated to form single embedding, and a classification layer is used to obtain the logits $H_{global}$ of all tree nodes. The logits are then split by the number of classes in each layer to obtain $H_{global}^{l}$.
The final class probabilities for each layer $P_{l}$ is then obtained by
$P_{l}=softmax(H_{local}^{l}+H_{global}^{l})$
Appendix B Implementation Details
The traditional single-turn model, the retriever, and the concatenation decision model used are using backbone initialized with $\Phi_{XLMR}$, a multi-lingual domain specific XLM-RoBERTa-base model continued to be pre-trained with contrastive learning. We use AdamW to finetune the backbone and all other modules with a learning rate of 5e-6 and 1e-3, respectively. In LARA, the LLM used is vicuna-13b-v1.5 on Hugging Face with 13B parameters. All test are run on a single Nvidia V100 GPU card with a 32GB of GPU memory. The number of demonstrations $K$ retrieved for each intent is set at 10 in this experiment. We also experimented with numbers below 10, the lower the number, the lower the performance. 10 is the highest number we can fit due to GPU memory constraint. Due to this constraint as well, the total number of tokens the in-context learning demonstrations can make up to are limited to 2300 tokens. If exceeded, the number of demonstrations in each candidate intent are pruned equally starting with the ones with the lowest cosine similarity scores to $q_{all}$. During inference time, if the generated intent does not match any of the provided options, the intent of $\mathcal{M}_{c}$ on $q_{n}$ will be considered as the final result.
B.1 Dataset Details
Table 4 shows the number of samples in each dataset. We have the single-turn training data available in abundance over the course of business operations after years. These single-turn samples will serve as the demonstration pool for in-context learning. To evaluate the effectiveness of our methods, we also have the CS teams to manually annotate some real multi-turn online sessions to serve as the test set. Each session queries $\mathcal{Q}$ will only have the last query $q_{n}$ labelled.
Market Lang. Intents Train(ST) Test(MT)
@@ B.1 Dataset Details
Table 4: The major languages, number of intents, and the number of samples in each market for Single Turn (ST) and Multi-Turn (MT).
Appendix C Prompt Demonstration
C.1 Prompt for ICL ($\mathcal{P}$)
To fit the width of this paper, we use SQ to represent Similar Question.
Prompt for ICL ($\mathcal{P}$)
# Task Description
A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the users questions. USER: Determine the intent for the targetted message from the examples, you must use the context in the history messages to arrive at the best answer.
# Examples
[Content] SQ_1 [Intent] Intent_name_1
[Content] SQ_2 [Intent] Intent_name_2
[Content] SQ_3 [Intent] Intent_name_3
# Note
DO NOT create new intent on your own, you must strictly use the intents in the examples.
DO NOT provide any explanation.
Output ONLY ONE intent for the targgetted message.
Consider the context from previous messages if the targetted message is unclear.
# Context
message 1: Users query
message 2: Users query with Entity
[Content] Last users query
# Output
ASSISTANT: [Intent] <$Model$ $generated$ $Intent\_name$>
C.2 Prompt for ICL ($\mathcal{P}_{symbolic}$)
In $\mathcal{P}_{symbolic}$ the original label name $l$ of each intent in $\mathcal{E}_{symbolic}$ are replaced with single-token symbols, e.g. ‘A’, ‘B’, …, which bear no meaning to the intents they represented. Explanation will be made in the instruction prompt $\mathcal{T}_{symbolic}$ to link the symbols back to their original intent label $y_{j}$, and the model is instructed to generated the symbols instead of full label names.
Prompt for ICL ($\mathcal{P}_{symbolic}$)
# Task Description
Content is Same as $\mathcal{P}$
# Examples
[Content] SQ_1 [Intent] A
[Content] SQ_1 [Intent] B
<$omitted$>
[Content] SQ_1 [Intent] B
# Intent options
A is Intent_name_1
B is Intent_name_2
# Note
Content is Same as $\mathcal{P}$
# Context
Format is same as $\mathcal{P}$
# Output
ASSISTANT: [Intent]
C.3 $\mathcal{P}_{prepend}$
In $\mathcal{P}_{prepend}$, representative symbols for each intent will be prepend to the original label name $l$, such that they are separated by an extra character as boundary, e.g. label “logistics$>$how long will it take to receive order?" will be represented as “A$>$logistics$>$how long will it take to receive order?". Note that the instruction prompt $\mathcal{T}$ remains the same, the trick is to limit the model generation token count to 1 on API level.
Prompt for ICL ($\mathcal{P}_{prepend}$)
# Task Description
Content is Same as $\mathcal{P}$
# Examples
[Content] SQ_1 [Intent] A>Intent_name_1
[Content] SQ_2 [Intent] B>Intent_name_2
<$omitted$>
[Content] SQ_3 [Intent] B>Intent_name_2
# Note
Content is Same as $\mathcal{P}$
# Context
Format is same as $\mathcal{P}$
# Output
ASSISTANT: [Intent] B
C.4 $\mathcal{P}_{formatted}$
Prompt for ICL ($\mathcal{P}_{formatted}$)
# Task Description
Content is Same as $\mathcal{P}$
# Examples
Format is same as $\mathcal{P}_{prepend}$
# Note
Content is Same as $\mathcal{P}_{prepend}$
# Context
[History msg 1] Query
[History msg 2] Query with Entity
[Content] that is the order id
# Output
ASSISTANT: [Intent]
Appendix D More Light-weight Deployment Method
The multi-component architecture can be complicated to implement for real-time systems. Alternatively, this method can be used offline as a multi-turn data pseudo-labeling tool to train a classification model. The training method will be the same as the $\mathcal{M}_{c}$ classifier in the paper, just with pseudo-labeled multi-turn data added to the original data with only single-turn samples. We also did experiment to ensure the quality of the model trained pseudo-labelled data.
The prompt used for the experiment is $\mathcal{P}_{formatted}$. Since this is not a real-time task, and we don’t need to care about the pipeline response time, we also did self-consistency checking on the LLM outputs to ensure the quality of pseudo-labels. For this checking, the in-context learning part is run three times per sample, with the in-context examples sorted in three fashions according to their scores: ascending, descending, and random. 70k of online chat logs are sampled for pseudo-labelling, and only those having consistent labels after 3 runs will be kept for training. Around 12% of the data will yield inconsistent results and be discarded. We validated that doing self-consistency this way can improve the average accuracy by 4.48% (from 64.64% to 69.12%), and thus the quality of the pseudo-label.
Moreover, the classifier trained with the high-quality multi-turn data generated by our pipeline can achieve better overall performance than the best original proposed method by 1.89% (64.64% vs 66.53%). This is all while cutting the deployment cost to just one classical classification model, but with the trade off of offline training time.